OpenStudy (anonymous):

One number is 8 less then a second number. Twice the first is 12 more than 4 times the second. Find the numbers. Value of the first number=? Value of the second number=?

OpenStudy (texaschic101):

x = first number y = second number y = x - 8 2x = 4y + 12 now we sub x - 8 in for y 2x = 4(x - 8) + 12 2x = 4x - 32 + 12 2x - 4x = -32 + 12 -2x = - 20 x = 10 y = x - 8 y = 10 - 8 y = 2 check.. 2x = 4y + 12 2(10) = 4(2) + 12 20 = 8 + 12 20 = 20 (correct) first number (x) = 10 second number (y) = 2

OpenStudy (anonymous):

Thank you!!!